### v1

@@ Abstract
Adding additional control to pretrained diffusion models has become an increasingly popular research area, with extensive applications in computer vision, reinforcement learning, and AI for science. Recently, several studies have proposed training-free diffusion guidance by using off-the-shelf networks pretrained on clean images. This approach enables zero-shot conditional generation for universal control formats, which appears to offer a free lunch in diffusion guidance. In this paper, we aim to develop a deeper understanding of the operational mechanisms and fundamental limitations of training-free guidance. We offer a theoretical analysis that supports training-free guidance from the perspective of optimization, distinguishing it from classifier-based (or classifier-free) guidance. To elucidate their drawbacks, we theoretically demonstrate that training-free methods are more susceptible to adversarial gradients and exhibit slower convergence rates compared to classifier guidance. We then introduce a collection of techniques designed to overcome the limitations, accompanied by theoretical rationale and empirical evidence. Our experiments in image and motion generation confirm the efficacy of these techniques.
@@ 1.1 Motivations
One critical area of research in the field of diffusion models involves enhancing controllability, such as pose manipulation in image diffusion [45], modulation of quantum properties in molecule diffusion [16], and direction of goal-oriented actions in RL diffusion [19]. The predominant techniques for exerting control over diffusion models include classifier guidance and classifier-free guidance. Classifier guidance involves training a time-dependent classifier to map a noisy image, denoted as $\bm{x}_{t}$, to a specific condition $\bm{y}$, and then employing the classifier’s gradient to influence each step of the diffusion process [7]. Conversely, classifier-free guidance bypasses the need for a classifier by training an additional diffusion model conditioned on $\bm{y}$ [15]. However, both approaches necessitate extra training to integrate the conditions. Moreover, their efficacy is often constrained when the data-condition pairs are limited and typically lacks the capability for zero-shot generalization to novel conditions.
@@
Recently, several studies [2, 44, 33] have introduced training-free guidance that builds upon the concept of classifier guidance. These models eschew the need for training a classifier on noisy images; instead, they estimate the clean image from its noisy counterpart using Tweedie’s formula and then employ pre-trained networks, designed for clean images, to guide the diffusion process. Given that checkpoints for these networks pretrained on clean images are widely accessible online, this form of guidance can be executed in a zero-shot manner. A unique advantage of training-free guidance is that it can be applied to universal control formats, such as style, layout, and FaceID [2, 44, 33] without any additional training efforts. Furthermore, these algorithms have been successfully applied to offline reinforcement learning, enabling agents to achieve novel goals not previously encountered during training [40]. Despite the encouraging empirical results, the fundamental principles contributing to the success of these methods have yet to be fully elucidated. From a theoretical perspective, it is intriguing to understand how and when these methods succeed or fail. From an empirical standpoint, it is crucial to develop algorithms that can address and overcome these limitations.
This paper seeks to deepen the understanding of training-free guidance by examining its operational mechanisms and inherent limitations. From an optimization standpoint, we demonstrate that training-free guidance has a two-stage convergence, where the second stage enjoys a linear convergence. This is a distinctive feature that is not shared by classifier guidance or classifier-free guidance methods. However, we identify and provide proof that training-free guidance is prone to adversarial gradient issues and slower convergence rates due to the diminished smoothness of the guidance network, when compared to classifier guidance. To address these challenges, we introduce a suite of techniques that have been theoretically and empirically validated for their efficacy. Specifically, our major contributions can be summarized as follows:
  • Mechanisms (How it works): We offer theoretical proof, from an optimization perspective, that training-free guidance is guaranteed to minimize the loss of the guidance network. This characteristic distinguishes them from approaches that rely on training-based guidance.
@@
  • Limitations (When it does not work): We theoretically identify the susceptibility of training-free guidance to adversarial gradient issues and slower convergence rates. We attribute these challenges to a decrease in the smoothness of the guidance network in contrast to the classifier guidance.
@@
  • Addressing the limitations: We introduce a set of enhancement techniques aimed at mitigating the identified limitations. The efficacy of these methods is empirically confirmed across various diffusion models (e.g., image diffusion and motion diffusion) and under multiple conditions (e.g., segmentation, sketch, text, and object avoidance). The code is available at https://github.com/BIGKnight/Understanding-Training-free-Diffusion-Guidance.
For more related fields and the impact of this paper in these fields, please refer to Appendix 7.
@@ 2.1 Diffusion Models
Diffusion models are characterized by forward and reverse processes. The forward process, occurring over a time interval from 00 to $T$, incrementally transforms an image into Gaussian noise. On the contrary, the reverse process, from $T$ back to 00, reconstructs the image from the noise. Let $\bm{x}_{t}$ represent the state of the data point at time $t$; the forward process systematically introduces noise to the data by following a predefined noise schedule given by
$\bm{x}_{t}=\sqrt{\alpha_{t}}\bm{x}_{0}+\sigma_{t}\bm{\epsilon}_{t},$
where $\alpha_{t}\in[0,1]$ is monotonically decreasing with $t$, $\sigma_{t}=\sqrt{1-\alpha_{t}}$, and $\bm{\epsilon}_{t}\sim\mathcal{N}(0,\bm{I})$ is random noise. Diffusion models use a neural network to learn the noise at each step:
@@
2.2 Training-Free Diffusion Guidance
@@
Classifier guidance corresponds to the energy function in (4) when the loss function employed is cross-entropy. This direct relationship does not hold when alternative loss functions, such as the $\ell_{2}$ distance [22], are utilized. Nonetheless, empirical research [20] indicates that classifier guidance remains effective across a diverse array of applications, including but not limited to text guidance and image similarity guidance.
Training-free guidance [2, 44, 6] puts the expectation in (4) inside the loss function:
$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t}):=$ $\nabla_{\bm{x}_{t}}\log\left[\exp(-\ell(f_{\phi}(\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}(\bm{x}_{0})),\bm{y})\right]$ (5)
=$\overset{(a)}{=}$ $-\nabla_{\bm{x}_{t}}\ell\left[f_{\phi}\left(\frac{\bm{x}_{t}-\sigma_{t}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)}{\sqrt{\alpha_{t}}}\right),\bm{y}\right],$
where (a) uses Tweedie’s formula $\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}(\bm{x}_{0})=\frac{\bm{x}_{t}-\sigma_{t}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)}{\sqrt{\alpha_{t}}}$. Leveraging this approximation permits the use of a pre-trained off-the-shelf network designed for processing clean data. The gradient of the last term in the energy function is obtained via backpropagation through both the guidance network and the diffusion backbone.
@@ 3 Analysis of Training-Free Guidance
3.1 Mechanisms of Training-free Guidance
3.1.1 The difficulty of approximating $p_{t}(\bm{y}|\bm{x}_{t})$ in high-dimensions.
From a probabilistic standpoint, it is observed that [33]
@@ 3.1 Mechanisms of Training-free Guidance
where $q(\bm{x}_{0}|\bm{x}_{t})=\mathcal{N}(\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}(\bm{x}_{0}),r_{t}^{2}\bm{I})$ and $r_{t}$ is a tunable parameter [33]. By taking $n=1$ and $r_{t}=0$, (6) is reduced to the guidance term in (5). As demonstrated in [33], the approximation is effective for one-dimensional distribution. However, the approximation denoted by (a) does not extend to high-dimensional data, such as images. This is due to the well-known high-dimensional probability phenomenon where $q(\bm{x}_{0}|\bm{x}_{t})$ tends to concentrate on a spherical shell centered at $\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}(\bm{x}_{0})$ with radius $r_{t}$, as discussed in Chapter 3.1 of [39]. Since the spherical shell represents a low-dimensional manifold with zero measure in the high-dimensional space, there is a significant likelihood that the supports of $p(\bm{x}_{0}|\bm{x}_{t})$ and $q(\bm{x}_{0}|\bm{x}_{t})$ do not overlap, rendering the approximation (a) ineffective.
3.1.2 Training-free guidance minimizes the loss from an optimization perspective.
Given the difficulty of analyzing the training-free guidance from a probability perspective, we turn to an optimization perspective for clarity. Denote $\hat{\bm{x}}_{0}=\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}(\bm{x}_{0})$, the gradient guidance term in (5) can be written as the following:
where the last equality follows the second-order Miyasawa relationship $\text{Cov}[\bm{x}_{0}|\bm{x}_{t}]=\sigma_{t}^{2}(\bm{I}+\sigma_{t}^{2}\nabla^{2}\log p_{t}(\bm{x}_{t}))$ [26]. Based on (7), we show that the guidance steps decrease the loss function associated with the clean classifier $f_{\phi}$, which is shown in the next proposition.
Proposition 1
Assume that the guidance loss function $\ell(f_{\phi}(\bm{x}_{0}),\bm{y})$ is $\mu$-PL (defined in Defintion 2 in appendix) and $L_{f}$-Lipschitz with respect to clean images $\bm{x}_{0}$, and the score function $\nabla\log p_{t}(\bm{x}_{t})$ is $L_{p}$-Lipschitz (defined in Defintion 1 in appendix) with respect to noisy image $\bm{x}_{t}$. Then the following conditions hold:
  1. 1.
[…]
  2. 2.
    Denote $\kappa_{1}=\frac{\mu\lambda_{\min}^{2}}{L_{f}(1+L_{p})\sqrt{\alpha_{t}}\sigma_{t}^{4}}$, After one gradient step $\hat{\bm{x}}_{t}=\bm{x}_{t}-\eta\nabla_{\bm{x}_{t}}\ell_{t}(\bm{x}_{t}),\eta=\frac{\sqrt{\alpha_{t}}}{L_{f}(1+L_{p})}$, we have $\ell_{t}(\hat{\bm{x}}_{t})\leq(1-\kappa_{1})\ell_{t}(\bm{x}_{t})$;
  3. 3.
    Consider a diffusion process that adheres to a bounded change in the objective function such that for any diffusion step, i.e., $\ell_{t-1}(\bm{x}_{t-1})\leq\frac{\ell_{t}(\hat{\bm{x}}_{t})}{(1-\kappa_{2})}$ for some $\kappa_{2}<\kappa_{1}$, then the objective function converges at a linear rate, i.e., $\ell_{t-1}(\bm{x}_{t-1})\leq\frac{1-\kappa_{1}}{1-\kappa_{2}}\ell_{t}(\bm{x}_{t})$.
@@
The proof is given in Appendix 8.2. The Lipschitz continuity and PL conditions are basic assumptions in optimization, and it has been shown that neural networks can locally satisfy these conditions [5]. Our next focus is on scenarios where the assumption in the third condition is met. Specifically, for large values of diffusion time $t$, the diffusion trajectory can exhibit substantial deviations between adjacent steps, leading to an increased value of $\kappa_{2}$ in the third condition. This phenomenon is less pronounced for smaller $t$. Consequently, the theorem leads us to anticipate a two-phase convergence pattern in training-free guidance: an initial phase where the loss remains relatively high and oscillates for larger $t$, followed by a swift decline (linear convergence) towards lower loss values as $t$ decreases. In Figure 1, we use ResNet-50 trained on clean images to guide ImageNet pre-trained diffusion models. The loss value at each diffusion step is plotted. As a reference, we choose $100$ images from the class “indigo bird” in ImageNet training set and compute the loss value, which is referred as “Correct Image Loss” in the figure. The two-stage convergence is verified by experiments. Remarkably, from theory and experiments, we see that training-free guidance can achieve low loss levels even in instances of guidance failure (Figure 0(b)). This is attributed to the influence of adversarial gradients, as discussed in the next subsection.
@@ 3.2 Limitations of Training-free Guidance
3.2.1 Training-free guidance is more sensitive to the adversarial gradient.
Adversarial gradient is a significant challenge for neural networks, which refers to minimal perturbations deliberately applied to inputs that can induce disproportionate alterations in the model’s output [34]. The resilience of a model to adversarial gradients is often analyzed through the lens of its Lipschitz constant [31]. If the model has a lower Lipschitz constant, then it is less sensitive to the input perturbations and thus is more robust. In classifier guidance, time-dependent classifiers are trained on noise-augmented images. Our finding is that adding Gaussian noise transforms non-Lipschitz classifiers into Lipschitz ones. This transition mitigates the adversarial gradient challenge by inherently enhancing the model’s robustness to such perturbations, as shown in the next proposition.
Proposition 2
@@
The proof is given in Appendix 8.3. We present visualizations of the accumulated gradients for both the time-dependent and off-the-shelf time-independent classifiers corresponding to different classes in Figure 1(b) and Figure 1(c), respectively. These visualizations are generated by initializing an image with a random background and computing $1000$ gradient steps for each classifier. For the time-dependent classifier, the input time for the $t$-th gradient step is $1000-t$. The images are generated purely by the classifier gradients without diffusion involved. For comparative analysis, we include the accumulated gradient of an adversarially robust classifier [30], as shown in Figure 1(a), which has been specifically trained to resist adversarial gradients. The resulting plots reveal a stark contrast: the gradient of the time-dependent classifier visually resembles the target image, whereas the gradient of the time-independent classifier does not exhibit such recognizability. This observation suggests that off-the-shelf time-independent classifiers are prone to generating adversarial gradients compared to the time-dependent classifier used in classifier guidance. It could potentially misdirect the guidance process. In contrast to yielding a direction that meaningfully minimizes the loss, the adversarial gradient primarily serves to minimize the loss in a manner that is not necessarily aligned with the intended guidance direction.
3.2.2 Training-free guidance slows down the convergence of reverse ODE.
The efficiency of an algorithm in solving reverse ordinary differential equations is often gauged by the number of non-linear function estimations (NFEs) required to achieve convergence. This metric is vital for algorithmic design, as it directly relates to computational cost and time efficiency [32]. In light of this, we explore the convergence rates associated with various guidance paradigms, beginning our analysis with a reverse ODE framework that incorporates a generic gradient guidance term. The formula is expressed as
@@
The proof is given in Appendix 8.4. Proposition 2 establishes that time-dependent classifiers exhibit superior gradient Lipschitz constants compared to their off-the-shelf time-independent counterparts. This disparity in smoothness attributes leads to a deceleration in convergence for training-free guidance methods, necessitating a greater number of NFEs to achieve the desired level of accuracy when compared to classifier guidance.

### v2

@@ Abstract
Adding additional control to pretrained diffusion models has become an increasingly popular research area, with extensive applications in computer vision, reinforcement learning, and AI for science. Recently, several studies have proposed training-free loss-based guidance by using off-the-shelf networks pretrained on clean images. This approach enables zero-shot conditional generation for universal control formats, which appears to offer a free lunch in diffusion guidance. In this paper, we aim to develop a deeper understanding of training-free guidance, as well as overcome its limitations. We offer a theoretical analysis that supports training-free guidance from the perspective of optimization, distinguishing it from classifier-based (or classifier-free) guidance. To elucidate their drawbacks, we theoretically demonstrate that training-free guidance is more susceptible to adversarial gradients and exhibits slower convergence rates compared to classifier guidance. We then introduce a collection of techniques designed to overcome the limitations, accompanied by theoretical rationale and empirical evidence. Our experiments in image and motion generation confirm the efficacy of these techniques.
@@ 1 Introduction
One critical area of research in the field of diffusion models involves enhancing controllability, such as pose manipulation in image diffusion [47], modulation of quantum properties in molecule diffusion [18], and direction of goal-oriented actions in RL diffusion [21]. The predominant techniques for exerting control over diffusion models include classifier guidance and classifier-free guidance. Classifier guidance involves training a time-dependent classifier to map a noisy image, denoted as $\bm{x}_{t}$, to a specific condition $\bm{y}$, and then employing the classifier’s gradient to influence each step of the diffusion process [8]. Conversely, classifier-free guidance bypasses the need for a classifier by training an additional diffusion model conditioned on $\bm{y}$ [16]. However, both approaches necessitate extra training to integrate the conditions. Moreover, their efficacy is often constrained when the data-condition pairs are limited and typically lack the capability for zero-shot generalization to novel conditions.
@@
Recently, several studies [2, 46, 35] have introduced training-free guidance that builds upon the concept of classifier guidance. These models eschew the need for training a classifier on noisy images; instead, they estimate the clean image from its noisy counterpart using Tweedie’s formula and then employ pre-trained networks, designed for clean images, to guide the diffusion process. Given that checkpoints for these networks pretrained on clean images are widely accessible online, this form of guidance can be executed in a zero-shot manner. A unique advantage of training-free guidance is that it can be applied to universal control formats, such as style, layout, and FaceID [2, 46, 35] without any additional training efforts. Furthermore, these algorithms have been successfully applied to offline reinforcement learning, enabling agents to achieve novel goals not previously encountered during training [42]. In contrast to classifier guidance and classifier-free guidance, it is proved in [25] that training-free guidance does not offer an approximation to the exact conditional energy. Therefore, from a theoretical perspective, it is intriguing to understand how and when these methods succeed or fail. From an empirical standpoint, it is crucial to develop algorithms that can address and overcome these limitations.
This paper seeks to deepen the understanding of training-free guidance by examining its mechanisms and inherent limitations, as well as overcoming these limitations. Specifically, our major contributions can be summarized as follows:
  • How does training-free guidance work? Although exact conditional energy is difficult to approximate in a training-free manner, from the optimization standpoint, we show that training-free guidance can effectively decrease the guidance loss function. The optimization perspective clarifies the mystery of why the guidance weights should be meticulously designed in relation to the guidance function and diffusion time, as observed in [46].
@@
  • When does training-free guidance not work? We theoretically identify the susceptibility of training-free guidance to adversarial gradient issues and slower convergence rates. We attribute these challenges to a decrease in the smoothness of the guidance network in contrast to the classifier guidance.
@@
  • Improving training-free guidance: We introduce random augmentation to alleviate the adversarial gradient and Polyak step size scheduling to improve convergence. The efficacy of these methods is empirically confirmed across various diffusion models (i.e., image diffusion and motion diffusion) and under multiple conditions (i.e., segmentation, sketch, text, object avoidance, and targeting) [^1] The code is available at
    https://github.com/BIGKnight/Understanding-Training-free-Diffusion-Guidance
@@ 2.1 Diffusion Models
Diffusion models are characterized by forward and reverse processes. The forward process, occurring over a time interval from 00 to $T$, incrementally transforms an image into Gaussian noise. On the contrary, the reverse process, from $T$ back to 00, reconstructs the image from the noise. Let $\bm{x}_{t}$ represent the state of the data point at time $t$; the forward process systematically introduces noise to the data by following a predefined noise schedule given by $\bm{x}_{t}=\sqrt{\alpha_{t}}\bm{x}_{0}+\sigma_{t}\bm{\epsilon}_{t}$, where $\alpha_{t}\in[0,1]$ is monotonically decreasing with $t$, $\sigma_{t}=\sqrt{1-\alpha_{t}}$, and $\bm{\epsilon}_{t}\sim\mathcal{N}(0,\bm{I})$ is random noise. Diffusion models use a neural network to learn the noise at each step:
@@
2.2 Diffusion Guidance
@@
Training-free loss-based guidance [2, 46, 6] puts the expectation in (4) inside the loss function:
$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t}):=$ $\nabla_{\bm{x}_{t}}\log\left[\exp(-\ell(f_{\phi}(\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}(\bm{x}_{0})),\bm{y})\right]\overset{(a)}{=}-\nabla_{\bm{x}_{t}}\ell\left[f_{\phi}\left(\frac{\bm{x}_{t}-\sigma_{t}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)}{\sqrt{\alpha_{t}}}\right),\bm{y}\right],$ (5)
where (a) uses Tweedie’s formula $\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}(\bm{x}_{0})=\frac{\bm{x}_{t}-\sigma_{t}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)}{\sqrt{\alpha_{t}}}$. Leveraging this formula permits the use of a pre-trained off-the-shelf network designed for processing clean data. The gradient of the last term in the energy function is obtained via backpropagation through both the guidance network and the diffusion backbone.
@@ 3 Analysis of Training-Free Guidance
3.1 How does Training-free Guidance Work?
On the difficulty of approximating $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t})$ in high-dimensional space. Despite its intuitive, [25] has shown that training-free guidance in (5) does not offer an approximation to the true energy in (4). The authors of [35] considers to directly approximate (4) with Gaussian distribution:
@@ 3.1 How does Training-free Guidance Work?
where $q(\bm{x}_{0}|\bm{x}_{t})$is chosen as $\mathcal{N}(\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}(\bm{x}_{0}),r_{t}^{2}\bm{I})$ and $r_{t}$ is a tunable parameter. As demonstrated in [35], the approximation is effective for one-dimensional distribution. However, we find that the approximation denoted by (a) does not extend to high-dimensional data (e.g., images) if the surrogate distribution $q$ is sub-Gaussian. This is due to the well-known high-dimensional probability phenomenon [41] that if $q$ has sub-Gaussian coordinates (e.g., iid and bounded), then $q(\bm{x}_{0}|\bm{x}_{t})$ tends to concentrate on a spherical shell centered at $\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}(\bm{x}_{0})$ with radius $r_{t}$ (details are in Appendix C.1). Since the spherical shell represents a low-dimensional manifold with zero measure in the high-dimensional space, there is a significant likelihood that the supports of $p(\bm{x}_{0}|\bm{x}_{t})$ and $q(\bm{x}_{0}|\bm{x}_{t})$ do not overlap, rendering the approximation (a) ineffective.
Understanding training-free guidance from an optimization perspective. We instead analyze the training-free guidance from the optimization perspective. Intuitively, in each step, the gradient is taken and the objective function decreases. At the initial stage of the diffusion ($t$ is large), the diffusion trajectory can exhibit substantial deviations between adjacent steps and may increase the objective value. So the objective value will oscillate at the beginning. When $t$ is smaller, the change to the sample is more fine-grained, leading to a bounded change in the objective value. Therefore, the objective value is guaranteed to decrease when $t$ is small, as showing in the next proposition.
Proposition 3.1.
Assume that the guidance loss function $\ell(f_{\phi}(\bm{x}_{0}),\bm{y})$ is $\mu$-PL (defined in Defintion D.2 in appendix) and $L_{f}$-Lipschitz with respect to clean images $\bm{x}_{0}$, and the score function $\nabla\log p_{t}(\bm{x}_{t})$ is $L_{p}$-Lipschitz (defined in Defintion D.1 in appendix) with respect to noisy image $\bm{x}_{t}$. Denote $\lambda_{\min}$ as the minimum eigenvalue of the (semi)-definite matrix $\text{Cov}[\bm{x}_{0}|\bm{x}_{t}]$. Then the following conditions hold: (1) Consider the loss function $\ell_{t}(\bm{x}_{t})=\ell\left(f_{\phi}\left(\frac{\bm{x}_{t}+\sigma_{t}^{2}\nabla\log p_{t}(\bm{x}_{t})}{\sqrt{\alpha_{t}}}\right),\bm{y}\right)$ and denote $\kappa_{1}=\frac{\mu\lambda_{\min}^{2}}{L_{f}(1+L_{p})\sqrt{\alpha_{t}}\sigma_{t}^{4}}$, After one gradient step $\hat{\bm{x}}_{t}=\bm{x}_{t}-\eta\nabla_{\bm{x}_{t}}\ell_{t}(\bm{x}_{t}),\eta=\frac{\sqrt{\alpha_{t}}}{L_{f}(1+L_{p})}$, we have $\ell_{t}(\hat{\bm{x}}_{t})\leq(1-\kappa_{1})\ell_{t}(\bm{x}_{t})$; (2) Consider a diffusion process that adheres to a bounded change in the objective function such that for any diffusion step, i.e., $\ell_{t-1}(\bm{x}_{t-1})\leq\frac{\ell_{t}(\hat{\bm{x}}_{t})}{(1-\kappa_{2})}$ for some $\kappa_{2}<\kappa_{1}$, then the objective function converges at a linear rate, i.e., $\ell_{t-1}(\bm{x}_{t-1})\leq\frac{1-\kappa_{1}}{1-\kappa_{2}}\ell_{t}(\bm{x}_{t})$.
@@
The proof is given in Appendix D.2. The Lipschitz continuity and PL conditions are basic assumptions in optimization, and it has been shown that neural networks can locally satisfy these conditions [5]. In Figure 1, we use ResNet-50 trained on clean images to guide ImageNet pre-trained diffusion models. The loss value at each diffusion step is plotted. As a reference, we choose $100$ images from the class “indigo bird” in ImageNet training set and compute the loss value, which is referred as “Correct Image Loss” in the figure. The objective value oscillates when $t$ is large, followed by a swift decrease, which verifies our analysis. Furthermore, the optimization perspective resolves the mystery of why the guidance weights (i.e., the step size in optimization) should be carefully selected with respect to the guidance function and time $t$ [46]. In [46], most guidance weights $\eta$ are chosen to be proportional to $\sqrt{\alpha_{t}}$ and dependent on guidance network, which differs from the weights used in classifier guidance and precisely aligns with our findings in Proposition 3.1. Remarkably, from theory and experiments, we see that training-free guidance can achieve low loss levels even in instances of guidance failure (Figure 1(b)). This is attributed to the influence of adversarial gradients, as discussed in the next subsection.
@@ 3.2 Limitations of Training-free Guidance
Training-free guidance is more sensitive to the adversarial gradient. Adversarial gradient is a significant challenge for neural networks, which refers to minimal perturbations deliberately applied to inputs that can induce disproportionate alterations in the model’s output [36]. The resilience of a model to adversarial gradients is often analyzed through the lens of its Lipschitz constant [32]. If the model has a lower Lipschitz constant, then the output is less sensitive to the input perturbations and thus is more robust. In classifier guidance, time-dependent classifiers are trained on noise-augmented images. Our finding is that adding Gaussian noise transforms non-Lipschitz classifiers into Lipschitz ones. This transition mitigates the adversarial gradient challenge by inherently enhancing the model’s robustness to such perturbations, as shown in the next proposition.
Proposition 3.2.
@@
The proof is given in Appendix D.3. We present visualizations of the accumulated gradients for both the time-dependent and off-the-shelf time-independent classifiers corresponding to different classes in Figure 2(b) and Figure 2(c), respectively. These visualizations are generated by initializing an image with a random background and computing $1000$ gradient steps for each classifier. For the time-dependent classifier, the input time for the $t$-th gradient step is $1000-t$. The images are generated purely by the classifier gradients without diffusion involved. For comparative analysis, we include the accumulated gradient of an adversarially robust classifier [33], as shown in Figure 2(a), which has been specifically trained to resist adversarial gradients. The resulting plots reveal a stark contrast: the gradient of the time-dependent classifier visually resembles the target image, whereas the gradient of the time-independent classifier does not exhibit such recognizability. This observation suggests that off-the-shelf time-independent classifiers are prone to generating adversarial gradients compared to the time-dependent classifier used in classifier guidance. In contrast to yielding a direction that meaningfully minimizes the loss, the adversarial gradient primarily serves to minimize the loss in a manner that is not necessarily aligned with the intended guidance direction.
Training-free guidance slows down the convergence of reverse ODE. The efficiency of an algorithm in solving reverse ordinary differential equations is often gauged by the number of non-linear function estimations (NFEs) required to achieve convergence. This metric is vital for algorithmic design, as it directly relates to computational cost and time efficiency [34]. In light of this, we explore the convergence rates associated with various guidance paradigms, beginning our analysis with a reverse ODE framework that incorporates a generic gradient guidance term. The formula is expressed as
@@
The proof is given in Appendix D.4. Proposition 3.2 establishes that time-dependent classifiers exhibit superior gradient Lipschitz constants compared to their off-the-shelf time-independent counterparts. This disparity in smoothness slows down the convergence for training-free guidance methods, necessitating a greater number of NFEs to achieve the desired level of accuracy when compared to classifier guidance.